# Dynamics of defects and interfaces for interacting quantum hard disks

### v1

@@ Abstract
Defects and interfaces are essential to understand the properties of matter. However, studying their dynamics in the quantum regime remains a challenge in particular concerning the regime of two spatial dimensions. Recently, it has been shown that a quantum counterpart of the hard-disk problem on a lattice yields defects and interfaces, which are stable just due to quantum effects while they delocalize and dissolve classically. Here, we study in more detail the properties of defects and interfaces in this quantum hard-disk problem with a particular emphasis on the stability of these quantum effects upon including perturbations. Specifically, we introduce short-range soft-core interactions between the hard disks. From both analytical arguments and numerical simulations we find that large classes of defects and interfaces remain stable even under such perturbations suggesting that the quantum nature of the dynamics exhibits a large range of robustness. Our findings demonstrate the stability and non-classical behavior of quantum interface dynamics, offering insights into the dynamics of two-dimensional quantum matter and establishing the quantum hard-disk model as a platform for studying unconventional constrained quantum dynamics.
@@ I Introduction
Interfaces in 2D quantum systems can exhibit rich dynamics driven by quantum effects. This includes for instance phase transitions decoupled from the bulk behavior – the roughening transition, where interfaces change from smooth to rugged configurations [1, 2]. The dynamic properties of these interfaces are equally complex, featuring non-trivial relaxation and potential non-ergodic behavior indicative of quantum many-body scars and Hilbert space fragmentation [3, 4, 5]. However, understanding real-time evolution in 2D quantum systems remains a significant theoretical challenge due to the complexity of simulating large, interacting many-body systems. Constrained quantum matter provides a particularly promising class of systems, as due to the constraints the effective Hilbert space explored during the dynamics remains limited. Recent advances in quantum simulators have brought the real-time dynamics of such systems within experimental reach [6, 7, 8, 9, 10, 11, 12, 13, 14]. In particular, Rydberg atomic systems are particularly promising in this context, as the so-called Rydberg blockade mechanism naturally implements constrained configurations [10].
In our previous work, we explored the dynamics of the quantum hard disk model [15], a lattice-based quantum analog of the classical hard disk model, where the constraint is induced by each particle carrying an excluded volume in the form of a disk (see Fig. 1(A)). While, the static equilibrium properties of the quantum model turned out to be exactly identical to the classical case with a symmetry-broken crystalline phase at high particle densities, the dynamics has revealed genuine quantum effects. The key observation in this previous work has been that there exists a large number of defects and interfaces of this crystalline phase which remain stable and nonergodic only in the quantum world. As a consequence, the quantum hard disk model on a lattice exhibits nonergodic behavior even in 2D due to the effect of quantum interference. A one-dimensional analogue of quantum hard rods has in the meantime also been realized experimentally [6].
While our previous results established the existence of stable quantum interfaces, one key open question has remained: to what extent do these structures persist when perturbations and deformations of the quantum hard-disk model are introduced? To address this, we extend our analysis by introducing additional short-range soft-core interactions and investigating their impact on defect and interface stability.
We find that such additional soft-core interactions give rise to a kind of dynamical heterogeneity, where distinct initial defect or interface configurations exhibit widely varying relaxation dynamics. Remarkably, we observe that some defects remain capable of retaining memory of the initial state for arbitrarily long times, a robustness that stems from the persistence of quantum many-body scars even in the presence of interactions. Meanwhile, other configurations that were exhibiting nonergodic behavior in the original quantum hard-disk model now display long-lived plateaus before eventually thermalizing. We also observe that for certain initial configurations, the lattice can exhibit phase coexistence, where portions become ergodic while other areas retain a crystalline structure. Our findings establish the quantum hard-disk model as a robust platform for studying constrained quantum interface dynamics, with implications for quantum simulations of non-ergodic behavior.
@@
This paper is organized as follows: In Section II, we introduce our model and its key components. Section III provides a concise summary of our previous findings on Hilbert space fragmentation, establishing the necessary context for our present study. In Section IV, we conduct a detailed analysis of the dynamics of both point-like and interface-like defects. Finally, in section V, we conclude by summarizing our results and discussing potential directions for future research.
@@
Figure 1: Defect dynamics on the square lattice with interactions. (A) Schematic representation of the interacting quantum hard disk model on a square lattice. (B) The Hilbert space fragmentation as a function of particle density $\eta$. Here, $\eta_{1}=1/L$, $\eta_{2}=\frac{1}{2}-\frac{\lceil L/2\rceil}{L^{2}}$, and $\eta_{\textrm{max}}=\frac{1}{2}-\frac{L(\textrm{mod}2)}{2L^{2}}$. $\eta^{*}$ denotes the threshold density for the weak-to-strong crossover region. (C) Long-time dynamics of on-site occupations at $Jt=10^{2}$, showing memory of the initial state even in the presence of interactions.
@@ II Model
$H=J\sum_{\langle i,j\rangle}P_{i}\left({a}_{i}^{\dagger}a_{j}+a_{j}^{\dagger}a_{i}\right)P_{j}+\lambda\sum_{\langle\langle i,j\rangle\rangle}n_{i}n_{j}\,\,.$ (1)
@@
The excluded volume effect due to the hard disks is implemented using projection operators $P_{i}=\prod_{j\ni|\vec{r}_{i}-\vec{r}_{j}|=1}(1-n_{j})/2,$ which prevent particles from occupying nearest-neighboring sites (see Fig. 1A for an illustration). Here, $n_{i}=a_{i}^{\dagger}a_{i}$ represents the occupation number operator. These constraints significantly reduce the computational basis size by excluding states that violate the hard-disk condition, making numerical simulations more efficient, see for instance Fig. 1 for a $20\times 20$ lattice. The terms $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ represent nearest-neighbor and next-to-nearest-neighbor pairs, respectively. Here, $\lambda$ denotes the strength of the next-to-nearest-neighbor interactions, while $J$ specifies the hopping strength, which is set to $J=1$ in our analysis. The model in Eq. (1) exhibits a $U(1)$-symmetry in terms of particle number conservation. In the following, we will therefore typically fix the particle density $\eta=M/L^{2}$ with $M$ denoting the total number of particles.
This model is, in principle, experimentally realizable, as demonstrated for the $1$D case in [6]. Of particular importance in this context are Rydberg atomic systems: the Rydberg blockade mechanism automatically enforces the excluded volume constraint by preventing the occupation of neighboring sites. The case of free hard disks ($\lambda=0$) was analyzed in detail in our previous work [15]. In this study, we extend our investigation by examining the robustness of these results in the presence of the additional soft-core interactions between the hard disks at $\lambda\not=0$.
@@ III Hilbert-Space fragmentation
The key structural element in our basis configurations that governs fragmentation in our model is what we term a snake—a fully filled diagonal within the configuration. The presence of a snake acts as a barrier, effectively dividing the lattice into two distinct regions and preventing particle movement between them, thus leading to fragmentation. At low particle densities, no snakes are present, and therefore, fragmentation does not occur. As the particle density $\eta$ increases to $1/L$, configurations can contain snakes. Beyond this threshold, the Hilbert space fragments into several smaller components, while a dominant large fragment persists, leading to weak fragmentation. As $\eta$ increases to $\eta>(1/2)-\lceil L/2\rceil/L^{2}$ (where $\lceil L/2\rceil$ denotes the ceiling function), each allowed configuration must include at least a snake or a fully frozen parallel diagonal. These frozen diagonals effectively separate the system into two distinct regions. Consequently, the Hilbert space becomes exclusively composed of small fragments, resulting in strong fragmentation.
@@ IV.2 Dynamics of Interfaces
Figure 4: Long-time average of the autocorrelation function $G(t)$ as a function of the interaction strength $\lambda$ corresponding to the initial configurations: (A) particles removed in the second row, (B) particles removed in the first row, and (C) particles removed along half of the diagonal.
@@
Figure 5: Quantum Many-Body Scars in the presence of interactions. The entanglement entropy of the eigenstates of the largest fragment is shown for different interaction strengths: (A) $\lambda=0$, representing free hard disks; (B) $\lambda=0.01$, showing weak interactions; (C) $\lambda=0.1$, illustrating moderate interactions; and (D) $\lambda=0.8$, highlighting strong interaction effects.
@@
The remarkable memory retention observed in the third scenario, in the presence of interactions, suggests a connection to quantum many-body scars, similar to those identified for free hard disks [15]. To further explore this phenomenon, we analyze the eigenstate properties using the bipartite entanglement entropy:
$S_{A}=-\operatorname{Tr}\{\operatorname{Tr}_{\bar{A}}\big{(}|\Psi\rangle\langle\Psi|\big{)}\,\textrm{ln}\left[\operatorname{Tr}_{\bar{A}}\big{(}|\Psi\rangle\langle\Psi|\big{)}\right]\}.$ (3)
Here, $|\Psi\rangle$ denotes an eigenstate, $A$ represents the subsystem chosen as the bottom half of the 2D lattice, and $\bar{A}$ is its complement.
@@
Additionally, to quantify how closely a given eigenstate resembles a specific many-body configuration, we employ the Edwards-Anderson (EA) order parameter $Q_{\textrm{EA}}=L^{-4}\sum_{i,j}|\langle\Psi|(2n_{i}-1)(2n_{j}-1)|\Psi\rangle|^{2}.$ We introduce a normalized version of the EA order parameter to ensure that it takes values between 0 and 1:
$Q=\frac{Q_{\textrm{EA}}-\left(2\eta-1\right)^{4}}{1-\left(2\eta-1\right)^{4}}.$ (4)
@@
With this normalization, $Q=1$ indicates that $|\Psi\rangle$ corresponds to a single configuration, while $Q=0$ implies that $|\Psi\rangle$ represents a uniform superposition over all basis states.
@@
In Fig. 5, we present $S_{A}$ for all eigenstates in the largest fragment in the weak fragmentation region as a function of energy density $E/L^{2}$ for various interaction strengths $\lambda$, at a fixed particle density $\eta=14/36$. The data points are further distinguished by the normalized EA order parameter $Q$, represented through a color gradient, providing additional insight into the structural characteristics of the eigenstates. To broaden this investigation beyond the largest fragment, Appendix B presents Fig. 7, which displays the same entanglement entropy versus energy density plot, now incorporating contributions from all fragments. This highlights the presence of additional scar states across different fragments.
@@
In Fig. 2(A), we illustrate the behavior for free hard disks. As expected from the Eigenstate Thermalization Hypothesis, a majority of the eigenstates follow the conventional ergodic paradigm. However, intriguingly, certain eigenstates exhibit low entanglement entropy $S_{A}$ and a high EA order parameter $Q$, even in the middle of the spectrum, deviating from ergodic expectations. These states are identified as quantum many-body scars, a signature of non-ergodic behavior.
In Figs. 2(B), (C), and (D), we present the behavior for the interacting case. Remarkably, even with interactions, quantum many-body scars persist. Although the number of such states decreases as the interaction strength $\lambda$ increases, a finite number of scars remain even for large $\lambda$.
@@
The presence of these scars directly impacts the dynamics, as initial states with a higher overlap with these eigenstates, such as the one shown in Fig. 2(C), retain their crystal structure indefinitely. This phenomenon is broadly accessible within the model and not fine-tuned, as indicated by the consistently high $Q$ values for a substantial number of these states. Moreover, despite all states belonging to the same fragment, the high $Q$ values observed for scar states indicate quantum interference effects as the underlying cause. This observation highlights the robustness of non-ergodic behavior in this model, demonstrating that it endures despite the introduction of interactions.
@@ V Conclusion and outlook
In this work, we studied the quantum hard disk model under the influence of interactions, focusing on its dynamical properties and the robustness of quantum many-body scars. Our results demonstrate that the defining features of the model, such as Hilbert-space fragmentation and non-ergodic dynamics, persist even with the inclusion of interactions. The introduction of interactions induces dynamical heterogeneity, with initial defect configurations exhibiting rapid memory loss, long-term decay, or indefinite retention of crystalline structure, depending on the interaction strength and initial configuration. Remarkably, quantum many-body scars remain robust in the presence of interactions, enabling certain initial states to retain memory over infinite timescales. These findings highlight the intricate interplay between constraints and interactions in constrained quantum systems and establish the quantum hard disk model as a versatile platform to explore non-ergodic dynamics and quantum many-body scars.
Future research could focus on developing a broader framework to understand the conditions under which Hilbert space fragmentation can coexist with quantum many-body scars. Additionally, it would be valuable to identify the specific properties of initial states that prevent them from thermalizing. An intriguing question for future exploration is whether our model’s unusual quantum behavior of defects and interfaces extends to other constrained quantum systems beyond hard disks. Given the fundamental role of interfaces in 2D quantum systems, understanding their stability and dynamics in constrained systems may lead us toward further understanding of non-ergodic behavior. Furthermore, an important next step would be to move beyond purely numerical approaches and develop an analytical understanding of the stability of interfaces. Investigating similar constrained models and identifying practical experimental setups to observe these effects would be crucial for advancing our understanding of non-ergodic quantum matter.

### v2

@@ Abstract
Defects and interfaces are essential to understand the properties of matter. However, studying their dynamics in the quantum regime remains a challenge in particular concerning the regime of two spatial dimensions. Recently, it has been shown that a quantum counterpart of the hard-disk problem on a lattice yields defects and interfaces, which are stable just due to quantum effects while they delocalize and dissolve in an analogous classical stochastic process. Here, we study in more detail the properties of defects and interfaces in this quantum hard-disk problem with a particular emphasis on the stability of these quantum effects upon including perturbations. Specifically, we introduce short-range soft-core interactions between the hard disks. From both analytical arguments and numerical simulations we find that large classes of defects and interfaces remain stable even under such perturbations suggesting that the quantum nature of the dynamics exhibits a large range of robustness. Our findings demonstrate the stability and non-classical behavior of quantum interface dynamics, offering insights into the dynamics of two-dimensional quantum matter and establishing the quantum hard-disk model as a platform for studying unconventional constrained quantum dynamics.
@@ I Introduction
Interfaces in two-dimensional quantum systems can exhibit remarkably rich dynamics driven by quantum interference and constraints. Among the most striking phenomena are interface-driven phase transitions that are decoupled from the bulk behavior, such as the roughening transition, where interfaces change from smooth to rugged configurations [1, 2]. These interfaces also display complex real-time dynamics, including non-trivial relaxation and signatures of non-ergodicity such as quantum many-body scars and Hilbert space fragmentation [3, 4, 5].
Recent works have further expanded this landscape by identifying a mechanism of ergodicity breaking in constrained quantum systems, namely quantum many-body caging  [6, 7, 8]. Here, quantum interference and local constraints cooperate to dynamically confine the system to a subset of configurations that form flat bands in the spectrum. These bands are associated with compact localized many-body eigenstates, which cannot hybridize with the rest of the Hilbert space due to destructive interference. As a result, certain initial states exhibit persistent memory and fail to thermalize, even in the absence of disorder or fine-tuning.
Despite these theoretical advances, simulating real-time evolution in 2D quantum many-body systems remains a formidable challenge. However, constrained quantum systems provide a promising route forward: the presence of local constraints significantly reduces the size of the dynamically accessible Hilbert space, making them amenable to both theoretical analysis and experimental simulation. In particular, Rydberg atom arrays, which naturally implement kinetic constraints through the Rydberg blockade mechanism, have emerged as an ideal platform for studying such dynamics [9, 10, 11, 12, 13, 14, 15, 16, 17].
In our previous work [18], we introduced and studied the quantum hard-disk model, a lattice-based quantum analog of the classical hard-disk system. Here, the exclusion constraint is realized via particles that carry a finite “excluded volume,” implemented as a hard-core radius around each particle (see Fig. 1(A)). While the model’s equilibrium properties matched the classical case—featuring a crystalline phase at high density—its dynamics revealed intrinsically quantum phenomena. Most notably, we identified a large number of defect and interface configurations that are dynamically stable only in the quantum model, owing to interference effects that prevent thermalization. As a result, the model exhibits nonergodic behavior even in two dimensions, defying conventional expectations for generic interacting systems. Recent developments have linked these nonergodic structures to the mechanism of quantum many-body caging, where constraints and interference stabilize long-lived atypical eigenstates [6].
While our prior work established the existence of stable quantum interfaces, a key open question remained: How robust are these structures under perturbations or deformations of the Hamiltonian? To answer this, we now extend our study by introducing additional short-range soft-core interactions, probing whether nonergodic behavior and defect stability persist beyond this limit. So this model is a compelling platform for studying nonergodicity, and they exhibit rich behavior. Moreover, a one-dimensional analog of the model, involving quantum hard rods, has now been realized experimentally [9], further motivating the study of this model.
We find that these interactions induce a form of dynamical heterogeneity, where different initial configurations exhibit markedly different relaxation dynamics. Some interface-like structures remain nonthermal for arbitrarily long times, retaining memory of their initial state. This robustness is tied to the persistence of quantum many-body cages, which survive even in the presence of interactions. Other configurations, by contrast, show long-lived plateaus followed by eventual thermalization, while some states exhibit coexistence of ergodic and nonergodic regions, revealing spatially resolved relaxation dynamics.
These findings highlight the quantum hard-disk model as a robust platform for exploring constrained quantum dynamics and nonergodicity in 2D, with relevance for current quantum simulation experiments and broader questions about the emergence of thermalization.
@@
This paper is organized as follows: In Section II, we introduce the model and the specific deformations we consider. Section III summarizes our previous results on constrained dynamics and Hilbert space fragmentation. Section IV presents a detailed analysis of the dynamical behavior of point-like and interface-like defects. Finally, Section V discusses implications and possible directions for future research.
@@
Figure 1: Defect dynamics on the square lattice with interactions. (A) Schematic representation of the interacting quantum hard disk model on a square lattice. The interactions are of next-to-nearest-neighbor type, acting only between particles located diagonally. There are no interactions between particles that are separated by two sites in the same row or column. (B) The Hilbert space fragmentation as a function of particle density $\eta$. Here, $\eta_{1}=1/L$, $\eta_{2}=\frac{1}{2}-\frac{\lceil L/2\rceil}{L^{2}}$, and $\eta_{\textrm{max}}=\frac{1}{2}-\frac{L(\textrm{mod}2)}{2L^{2}}$. $\eta^{*}$ denotes the threshold density for the weak-to-strong crossover region. (C) Long-time dynamics of on-site occupations at $Jt=10^{2}$, showing memory of the initial state even in the presence of interactions.
@@ II Model
$H=J\sum_{\langle i,j\rangle}P_{i}{a}_{i}^{\dagger}a_{j}P_{j}+P_{j}a_{j}^{\dagger}a_{i}P_{i}+\lambda\sum_{\langle\langle i,j\rangle\rangle}n_{i}n_{j}\,\,.$ (1)
@@
The excluded volume effect due to the hard disks is implemented using projection operators $P_{i}=\prod_{j\ni|\vec{r}_{i}-\vec{r}_{j}|=1}(1-n_{j}),$ which prevent particles from occupying nearest-neighboring sites (see Fig. 1A for an illustration). Here, $n_{i}=a_{i}^{\dagger}a_{i}$ represents the occupation number operator. These constraints significantly reduce the computational basis size by excluding states that violate the hard-disk condition, making numerical simulations more efficient, see for instance Fig. 1 for a $20\times 20$ lattice. The terms $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ represent nearest-neighbor and next-to-nearest-neighbor pairs, respectively. Here, $\lambda$ denotes the strength of the next-to-nearest-neighbor interactions, while $J$ specifies the hopping strength, which is set to $J=1$ in our analysis. The model in Eq. (1) exhibits a $U(1)$-symmetry in terms of particle number conservation. In the following, we will therefore typically fix the particle density $\eta=M/L^{2}$ with $M$ denoting the total number of particles.
This model is, in principle, experimentally realizable, as demonstrated for the 11D case in [9]. Of particular importance in this context are Rydberg atomic systems: the Rydberg blockade mechanism automatically enforces the excluded volume constraint by preventing the occupation of neighboring sites. The case of free hard disks ($\lambda=0$) was analyzed in detail in our previous work [18]. In this study, we extend our investigation by examining the robustness of these results in the presence of the additional soft-core interactions between the hard disks at $\lambda\not=0$.
@@ III Hilbert-Space fragmentation
The key structural element in our basis configurations that governs fragmentation in our model is what we term a snake—a fully filled main diagonal or its anti-diagonal within the configuration. The presence of a snake acts as a barrier, effectively dividing the lattice into two distinct regions and preventing particle movement between them, thus leading to fragmentation. At low particle densities, no snakes are present, and therefore, fragmentation does not occur. As the particle density $\eta$ increases to $1/L$, configurations can contain snakes. Beyond this threshold, the Hilbert space fragments into several smaller components, while a dominant large fragment persists, leading to weak fragmentation. As $\eta$ increases to $\eta>(1/2)-\lceil L/2\rceil/L^{2}$ (where $\lceil L/2\rceil$ denotes the ceiling function), each allowed configuration must include at least a snake or a fully frozen parallel diagonal. These frozen diagonals effectively separate the system into two distinct regions. Consequently, the Hilbert space becomes exclusively composed of small fragments, resulting in strong fragmentation.
@@ IV.2 Dynamics of Interfaces
Figure 4: Long-time average of the autocorrelation function $G(t)$ as a function of the interaction strength $\lambda$ corresponding to the initial configurations: (A) particles removed in the second row, (B) particles removed in the first row, and (C) particles removed along half of the diagonal. The time-window considered here is $Jt=10^{4}-10^{5}.$
@@
[…]
@@
The remarkable memory retention observed in the third scenario, in the presence of interactions, suggests a connection to atypical eigenstates, similar to those identified for free hard disks [18, 6]. To further explore this phenomenon, we analyze the eigenstate properties using the bipartite entanglement entropy:
[…]
Here, $|\Psi\rangle$ denotes an eigenstate, $A$ represents the subsystem chosen as the bottom half of the 2D lattice, and $\bar{A}$ is its complement. Our Hamiltonian possesses rotational and reflectional symmetries, which give rise to degeneracies. As a consequence, the entanglement entropy is not uniquely defined, since any superposition within a degenerate manifold is also an eigenstate with the same energy, leading to different entropy values depending on the chosen superposition. To eliminate this ambiguity, we work in a symmetry-resolved basis. Specifically, we apply the full dihedral symmetry group of order four on the $6\times 6$ lattice and diagonalize the Hamiltonian within the $A_{1}$ irreducible representation. This procedure enforces the lattice symmetries at the level of the Hilbert space itself and yields uniquely defined eigenstates.
@@
Additionally, to quantify how closely a given eigenstate resembles a specific many-body configuration, we employ the Edwards-Anderson (EA) order parameter $Q_{\textrm{EA}}=L^{-4}\sum_{i,j}|\langle\Psi|(2n_{i}-1)(2n_{j}-1)|\Psi\rangle|^{2}.$ We introduce a normalized Edwards–Anderson (EA) order parameter to ensure that it lies between 0 and 1:
$Q=\frac{Q_{\mathrm{EA}}-(2\eta-1)^{4}}{Q_{\mathrm{EA}}^{\mathrm{max}}-(2\eta-1)^{4}}.$ (4)
@@
Here, $Q_{\mathrm{EA}}^{\mathrm{max}}$ denotes the maximal value of $Q_{\mathrm{EA}}$ attainable within the $A_{1}$ irrep subspace. With this normalization, $Q=1$ indicates that $|\Psi\rangle$ corresponds to a single basis state of the subspace, whereas $Q=0$ signifies that $|\Psi\rangle$ is a uniform superposition over all basis states.
@@
In Fig. 5, we present $S_{A}$ for all eigenstates in the largest fragment in the weak fragmentation region as a function of energy density $E/L^{2}$ for various interaction strengths $\lambda$, at a fixed particle density $\eta=14/36$. The data points are further distinguished by the normalized EA order parameter $Q$, represented through a color gradient, providing additional insight into the structural characteristics of the eigenstates. To broaden this investigation beyond the largest fragment, Appendix B presents Fig. 7, which displays the same entanglement entropy versus energy density plot, now incorporating contributions from all fragments. This highlights the presence of additional non-ergodic states across different fragments.
@@
In Figs. 5(A), (B), and (C), as expected from the Eigenstate Thermalization Hypothesis, a majority of the eigenstates follow the conventional ergodic paradigm. However, intriguingly, certain eigenstates exhibit low entanglement entropy $S_{A}$ and a high EA order parameter $Q$, even in the middle of the spectrum, deviating from ergodic expectations. These states are identified in the quantum hard disk model as quantum many-body cages, a recently developed signature of non-ergodic behavior [6]. Remarkably, our results suggest that even with interactions, quantum many-body cages persist. Although the number of such states decreases as the interaction strength $\lambda$ increases, a finite number of cages remain even for large $\lambda$. Notice that the prominent towers of eigenstates visible in the spectrum are not exact degeneracies. These towers become exactly degenerate only at $\lambda=0$. Once interactions are introduced, the levels split and the degeneracy is lifted.
@@
The presence of these cages directly impacts the dynamics, as initial states with a higher overlap with these eigenstates, such as the one shown in Fig. 2(C), retain their crystal structure indefinitely. This phenomenon is broadly accessible within the model and not fine-tuned, as indicated by the consistently high $Q$ values for a substantial number of these states. Moreover, despite all states belonging to the same fragment, the high $Q$ values observed for cages states indicate quantum interference effects as the underlying cause. This observation highlights the robustness of non-ergodic behavior in this model, demonstrating that it endures despite the introduction of interactions.
@@ V Conclusion and outlook
In this work, we studied the quantum hard disk model under the influence of interactions, focusing on its dynamical properties and the robustness of quantum many-body cages. Our results demonstrate that the defining features of the model, such as Hilbert-space fragmentation and non-ergodic dynamics, persist even with the inclusion of interactions. The introduction of interactions induces dynamical heterogeneity, with initial defect configurations exhibiting rapid memory loss, long-term decay, or indefinite retention of crystalline structure, depending on the interaction strength and initial configuration. Remarkably, quantum many-body cages remain robust in the presence of interactions, enabling certain initial states to retain memory over infinite timescales. These findings highlight the intricate interplay between constraints and interactions in constrained quantum systems and establish the quantum hard disk model as a versatile platform to explore non-ergodic dynamics and quantum many-body cages. Future research could focus on developing a broader framework to understand the robustness of quantum many-body cages due to perturbation. Additionally, it would be valuable to identify the specific properties of initial states that prevent them from thermalizing. An intriguing question for future exploration is whether our model’s unusual quantum behavior of defects and interfaces extends to other constrained quantum systems beyond hard disks. Given the fundamental role of interfaces in 2D quantum systems, understanding their stability and dynamics in constrained systems may lead us toward further understanding of non-ergodic behavior. Furthermore, an important next step would be to move beyond purely numerical approaches and develop an analytical understanding of the stability of interfaces. Investigating similar constrained models and identifying practical experimental setups to observe these effects would be crucial for advancing our understanding of non-ergodic quantum matter.